OpenStudy (loser66):

What is the result? (-1)^(n^2/2) = 1if n is even what is the result if n is odd? Please, help

OpenStudy (loser66):

@dan815

OpenStudy (danjs):

try it

OpenStudy (loser66):

Can you?

OpenStudy (loser66):

This is the original problem: Find the radius convergence of \(\sum_{n=1}^\infty \dfrac{i^{n^2}}{n}\)

OpenStudy (loser66):

We need an argument for i^n^2

Parth (parthkohli):

To be fair the root of \(-1\) can be \(+i\) or \(-i\).

imqwerty (imqwerty):

n=even that thing is 1 n=odd then its i

imqwerty (imqwerty):

wait n=odd has got cases

Parth (parthkohli):

Exactly, we can't tell if it's \(i\) or \(-i\) right?

OpenStudy (loser66):

\(i^{n^2}= (i^2)^{n^2/2}=(-1)^{n^2/2}\)

imqwerty (imqwerty):

yea tru

OpenStudy (loser66):

ok, then , right?